In the presence of a magnetic field with a strength of 6.0x10^-3T, a 36m long wire carries a


current of 11.0A. Determine the amount of the magnetic force acting on the wire. The angle of


the line is 45 degrees

1
Expert's answer
2022-05-12T15:09:20-0400
"F=BIl\sin\theta\\
=6.0*10^{-3}*11.0*36*\sin45^\circ=1.6\:\rm N"